# A Mechanistic Interpretation of Syllogistic Reasoning in Auto-Regressive Language Models

### v2

@@ Abstract
Recent studies on logical reasoning in Language Models (LMs) have sparked a debate on whether they can learn systematic reasoning principles during pre-training or merely exploit superficial patterns in the training data. This paper presents a mechanistic interpretation of syllogistic reasoning in LMs to advance the understanding of internal dynamics. Specifically, we present a methodology for circuit discovery aimed at interpreting content-independent reasoning mechanisms. Through two distinct intervention methods, we uncover a sufficient and necessary circuit involving middle-term suppression that elucidates how LMs transfer information to derive valid conclusions from premises. Furthermore, we investigate how belief biases manifest in syllogistic reasoning, finding evidence of partial contamination from additional attention heads responsible for encoding commonsense and contextualized knowledge. Finally, we explore the generalization of the discovered mechanisms across various syllogistic schemes, model sizes and architectures, finding that the identified circuit is sufficient and necessary for the schemes on which the models achieve high downstream accuracy ($\geq$ 60%), and that the activation patterns apply to models of different families. Overall, our findings suggest that LMs indeed learn transferable content-independent reasoning mechanisms, but that, at the same time, such mechanisms do not involve generalizable and abstract logical primitives, being susceptible to contamination by the same world knowledge acquired during pre-training.
A Mechanistic Interpretation of Syllogistic Reasoning in
Auto-Regressive Language Models
Geonhee Kim1, Marco Valentino3, André Freitas1,2,3 1Department of Computer Science, University of Manchester, UK 2National Biomarker Centre (NBC), CRUK Manchester Institute, UK 3 Idiap Research Institute, Switzerland
@@ 1 Introduction
Transformer-based Language Models (LMs) have led to remarkable results across various natural language processing tasks (Radford et al., 2018, 2019; Brown et al., 2020; OpenAI et al., 2024; Jason et al., 2022; Bubeck et al., 2023). This success has catalyzed research interest in systematically exploring the reasoning capabilities emerging during pre-training (Clark et al., 2020). Recent findings suggest that logical reasoning abilities may emerge in large-scale models (Rae et al., 2022; Kojima et al., 2022; Wei et al., 2022) or through transfer learning on specialized datasets (Betz et al., 2021). However, ongoing debates question whether these models apply systematic inference rules or reuse superficial patterns learned during pre-training (Talmor et al., 2020; Kassner et al., 2020; Wu et al., 2024). This controversy underscores the need for a deeper understanding of the low-level logical inference mechanisms in LMs (Rozanova et al., 2024, 2023, 2022).
To improve our understanding of the internal mechanisms, this paper focuses on Mechanistic Interpretability (Olah et al., 2020; Nanda et al., 2023), aiming to discover the core circuit responsible for syllogistic reasoning. In particular, we focus on categorical syllogisms with universal affirmative quantifiers (i.e., AAA-1, Barbara) motivated by two key factors. First, as observed in natural logic studies (MacCartney and Manning, 2007), this form of syllogistic reasoning is prevalent in everyday language. Therefore, it is likely that LMs are exposed to such reasoning schema during pre-training. Second, AAA-1 is a form of unconditionally valid syllogism independent of the premises’ truth condition (Holyoak and Morrison, 2005). This characteristic offers a deterministic and scalable task design, other than allowing us to investigate the disentanglement between reasoning and knowledge representation Wysocka et al. (2024); Bertolazzi et al. (2024); Lampinen et al. (2024).
Through mechanistic interpretability techniques such as Activation Patching (Meng et al., 2022) and embedding space analysis (i.e., Logit Lens) (Nostalgebraist, 2020; Geva et al., 2022; Dar et al., 2023), we investigate the following main research questions: RQ1: How is the content-independent syllogistic reasoning mechanism internalized in LMs during pre-training?; RQ2: Are content-independent mechanisms disentangled from specific world knowledge and belief biases?; RQ3: Does the core reasoning mechanism generalize across model sizes and syllogistic schemes?
@@
1. The circuit analysis reveals that LMs develop specific inference mechanisms during pre-training, finding evidence supporting a three-stage mechanism for syllogistic reasoning: (1) naive recitation of the first premise; (2) suppression of duplicated middle-term information; and (3) mediation towards the correct output through the interplay of attention heads.
@@
3. We found that the identified circuit is sufficient and necessary for all the syllogistic schemes in which the model achieves high downstream accuracy ($\geq$ 60%). This result suggests that LMs learn reasoning mechanisms that are transferable across different schemes.
4. The intervention results on models with different architectures and sizes (i.e., GPT-2 (Radford et al., 2019), Pythia (Biderman et al., 2023), Llama (Grattafiori et al., 2024), Qwen (Yang et al., 2024)) show similar suppression mechanism patterns and information flow. However, we found evidence that the contribution of attention heads becomes more complex with model sizes, further supporting the hypothesis of increasing contamination from external world knowledge.
@@ 2.2 Circuit Discovery
This technique involves modifying the activation of targeted components and observing the resulting changes. Our study primarily examines the activation of residual streams and multi-head self-attention at the sequence level to trace the term information flow. Given a masked syllogistic input $x=[\mathcal{P}_{1};\mathcal{P}_{2};\mathcal{C}\setminus\{p\}]$ as $(s,m_{1},m_{2},p)$, which can be read as “All $s$ are $m_{1}$. All $m_{2}$ are $p$. Therefore, all $s$ are [mask]”, and its correct completion $y=p$, the Activation Patching technique consists of the following steps:
@@
For the target input $(s,m_{1},m_{2},p)$, we record the baseline logit difference, $\delta_{+}(p,m)$ from the forward pass output of the model (Figure 2(1)).
@@
Figure 3: Results of the symbolic circuit analysis. (a) Attention output patching and (b) residual stream patching obtained via middle-term intervention. (c) OV circuit logit lens results for attention head $\mathbf{h}_{11.10}$, with input and output comprising 26 uppercase letters. (d) Residual stream patching obtained via all-term corruption. (e)-(h) Average attention weights across the batch: (e) weights for token position $[p]$, and (f)-(h) weights for the last token position. For clarity, the symbol ‘-’ on the y-axis denotes the averaged values for tokens appearing between terms.
2.3 Interventions
@@ 3 Empirical Evaluation
We select the LM for the circuit analysis based on a trade-off between model size and performance. To this end, we measure the accuracy of GPT-2 (Radford et al., 2019) with different sizes (117M, 345M, 762M, and 1.5B) on the syllogism completion task. Specifically, we aim to identify potential phase transition points where significant performance changes emerge (Jason et al., 2022; Kaplan et al., 2020). As shown in Figure 6(a), we observe a notable transition from small to medium sizes, where averaged performance in three different datasets increases by 439.06%. The shift is more apparent under the logit difference metric (See Section 3.8 and Figure 6(b)). Given these results, we select GPT-2 Medium for circuit discovery and subsequently test the compatibility with different model sizes and architectures (Section 3.7 and 3.8).
@@ 3.1 Transitive Reasoning Mechanisms
We further investigate how information from $[m_{2}]$ is moved to $[p]$, positing that this relates to the model’s internal reasoning mechanism. We notice that at position $[p]$, the model can observe the complete AAA-1 syllogistic structure (Figure 2(a)) with middle-term duplication. We hypothesize that this structural information for reasoning is collected at one position, given that information refinement occurs at the specific position such as last token (Hanna et al., 2024; Stolfo et al., 2023). To verify this, we trace the information flow of head $\mathbf{h}_{11.10}$ using path patching (Wang et al., 2023). Our results show that $\mathbf{h}_{11.10}$ operates based on several induction heads (Elhage et al., 2021) ($\mathbf{h}_{5.8}$, $\mathbf{h}_{6.1}$, $\mathbf{h}_{6.15}$ and $\mathbf{h}_{7.2}$) formed at $[m_{2}]$. These heads attend to the $[[m_{1}]+1]$ token due to the $m_{1}\equiv m_{2}$ conditioned matching operation, likely containing $m_{1}$-related information from previous token heads. We conclude, therefore, that $m_{1}$ and $m_{2}$ information are aggregated at position $[p]$. Additional details of the path patching results are available in the Appendix G.
Suppressive mechanism is revealed through Logit Lens.
@@ 3.2 Term-Related Information Flow
Given the findings that the information for prediction resides in term-specific positions, we trace which attention heads transfer the information from each term position ($[p]$, $[m_{1}]$ and $[m_{2}]$) to the last token. We call these heads mover heads as the existence of “positive or negative mover heads” that carry or suppress information from the specific token position to the last token position (Wang et al., 2023; García-Carrasco et al., 2024). To discover the source of information flow, we investigate the heads with high attention value-based patching score and identify 9 notable mover heads (see Appendix H for details of mover heads localization). By further investigation of attention patterns (Figure 3(f)–(h)), $\mathbf{h}_{14.14},\mathbf{h}_{15.14}$ and $\mathbf{h}_{18.12}$ are classified as positive copy heads attending highly to $[p]$, and $\mathbf{h}_{19.1}$ is a positive suppression head by attending to $[m_{2}]$ and $[s]$. $\mathbf{h}_{9.9}$, $\mathbf{h}_{11.1}$, $\mathbf{h}_{12.1}$, $\mathbf{h}_{17.2}$ and $\mathbf{h}_{23.10}$ are negative copy heads giving a strong attention to non-$[p]$.
@@
3.3 Summary
Overall, we found that the mechanisms to derive correct conclusion involve the following phases:
@@ 3.3 Summary
Token-specific information is propagated to the last token position. The process culminates in the prediction shift from ‘B’ to the correct token ‘C’.
Therefore, these results suggest that the discovered circuit is mainly characterized by an internal error correction mechanism.
3.4 Circuit Evaluation
@@
Figure 4: Circuit ablation results for (a) correctness and (b) robustness on the symbolic dataset. Solid lines show numeric perturbation-based performance, dotted lines indicate quantifier perturbation-based performance in (b), and the dashed line shows the baseline logit difference without knockouts.
Figure 5: Circuit ablation results for evaluating transferability on non-symbolic datasets: the belief-consistent (a) and the belief-inconsistent (b). The dashed line represents the baseline logit difference without knockouts.
@@
3.5 Circuit Transferability
The circuit is necessary for non-symbolic syllogistic arguments, yet not sufficient for belief-inconsistent ones.
We present ablation results for the two generated non-symbolic datasets: belief-consistent (Figure 5(a)) and belief-inconsistent (Figure 5(b)) (See Appendix I for the intervention results). Notably, the symbolic circuit proves necessary for both types of non-symbolic inputs, suggesting that the logic derived from symbolic syllogisms remains essential even when natural words are substituted. Regarding sufficiency, while belief-consistent data show significant performance recovery, we observe an inability to restore original performance on the belief-inconsistent set despite an increasing trend. These results indicate that belief biases encoded in different attention heads may play an important role.
@@ 3.6 Generalization to Syllogistic Schemes
Table 2: Generalizability across unconditionally valid syllogistic forms. Columns C1, C2, and C3 indicate whether conditions are met ($\checkmark$) or not (-): C1: Necessity, C2: Sufficiency, C3: Positive Logit Difference. ‘Accuracy (Acc)’ shows performance on the completion task for each syllogism.
@@
We further aim to understand whether the symbolic circuit is specific to the AAA-1 syllogism (Barbara). To this end, we extend our experiment to encompass all 15 unconditionally valid syllogisms (see Appendix A and J for details about the syllogistic schemes and more experimental results). To evaluate the transferability to all 15 schemes, we verify three main conditions, reported in Table 2: (C1) Necessity, (C2) Sufficiency, and (C3) Positive Logit Difference. We also measure the accuracy of the completion task for each syllogism. From Table 2, we observe that the circuit is sufficient and necessary for all the syllogistic schemes in which the model achieves high downstream accuracy ($\geq$ 60%), supporting the conclusion that the circuit includes components that are crucial for the emergence of syllogistic reasoning in general. Notably, all three conditions are satisfied for the affirmative syllogisms (AII-3, IAI-3, IAI-4), for which the model achieves an accuracy above 0.6.
3.7 Generalization to Different Sizes

### v3

# Reasoning Circuits in Language Models: A Mechanistic Interpretation of Syllogistic Inference

@@ Abstract
Recent studies on reasoning in language models (LMs) have sparked a debate on whether they can learn systematic inferential principles or merely exploit superficial patterns in the training data. To understand and uncover the mechanisms adopted for formal reasoning in LMs, this paper presents a mechanistic interpretation of syllogistic inference. Specifically, we present a methodology for circuit discovery aimed at interpreting content-independent and formal reasoning mechanisms. Through two distinct intervention methods, we uncover a sufficient and necessary circuit involving middle-term suppression that elucidates how LMs transfer information to derive valid conclusions from premises. Furthermore, we investigate how belief biases manifest in syllogistic inference, finding evidence of partial contamination from additional attention heads responsible for encoding commonsense and contextualized knowledge. Finally, we explore the generalization of the discovered mechanisms across various syllogistic schemes, model sizes and architectures. The identified circuit is sufficient and necessary for syllogistic schemes on which the models achieve high accuracy ($\geq$ 60%), with compatible activation patterns across models of different families. Overall, our findings suggest that LMs learn transferable content-independent reasoning mechanisms, but that, at the same time, such mechanisms do not involve generalizable and abstract logical primitives, being susceptible to contamination by the same world knowledge acquired during pre-training.
Reasoning Circuits in Language Models:
A Mechanistic Interpretation of Syllogistic Inference
Geonhee Kim1, Marco Valentino2,3, André Freitas1,2,4 1Department of Computer Science, University of Manchester, UK 2Idiap Research Institute, Switzerland 3School of Computer Science, University of Sheffield, UK 4National Biomarker Centre, CRUK-MI, University of Manchester, UK kimgeonhee317@gmail.com, ac4mv@sheffield.ac.uk, andre.freitas@idiap.ch
@@ 1 Introduction
Language models (LMs) have led to remarkable results across various natural language processing tasks (Radford et al., 2018, 2019; Brown et al., 2020; OpenAI, 2024; Jason et al., 2022; Bubeck et al., 2023). This success has catalyzed research interest in systematically exploring the reasoning capabilities emerging during pre-training (Clark et al., 2020). Recent findings suggest that logical and formal reasoning abilities may emerge in large-scale models (et al., 2022; Kojima et al., 2022; Wei et al., 2022) or through transfer learning on specialized datasets (Betz et al., 2021). However, ongoing debates question whether these models apply systematic inference rules or reuse superficial patterns learned during pre-training (Talmor et al., 2020; Kassner et al., 2020; Wu et al., 2024). This controversy underscores the need for a deeper understanding of the low-level logical inference mechanisms in LMs (Rozanova et al., 2024, 2023, 2022; Yanaka et al., 2020).
To improve our understanding of the internal mechanisms, this paper focuses on mechanistic interpretability (Olah et al., 2020; Nanda et al., 2023), aiming to discover the core circuit responsible for syllogistic reasoning. In particular, we focus on categorical syllogisms with universal affirmative quantifiers (i.e., AAA-1, Barbara) motivated by two key factors. First, as observed in natural logic studies (MacCartney and Manning, 2007), this form of syllogistic reasoning is prevalent in everyday language. Therefore, it is likely that LMs are exposed to such reasoning schema during pre-training. Second, AAA-1 is a form of unconditionally valid syllogism independent of the premises’ truth condition (Holyoak and Morrison, 2005). This characteristic offers a deterministic and scalable task design, other than allowing us to investigate the disentanglement between reasoning and knowledge representation Bertolazzi et al. (2024); Wysocka et al. (2025); Lampinen et al. (2024); Valentino et al. (2025).
Through mechanistic interpretability techniques such as Activation Patching (Meng et al., 2022) and embedding space analysis (i.e., Logit Lens) (Nostalgebraist, 2020; Geva et al., 2022; Dar et al., 2023), we investigate the following main research questions: RQ1: How is the content-independent syllogistic reasoning mechanism internalized in LMs during pre-training?; RQ2: Are content-independent mechanisms disentangled from specific world knowledge and belief biases?; RQ3: Does the core reasoning mechanism generalize across syllogistic schemes, different model sizes, and architectures?
@@
1. The circuit analysis reveals that LMs develop specific inference mechanisms during pre-training, finding evidence supporting a three-stage mechanism for syllogistic reasoning: (1) naive recitation of the first premise; (2) suppression of duplicated middle-term information; and (3) mediation towards the correct output through the interplay of mover attention heads (see Figure 1).
@@
3. We found that the identified circuit is sufficient and necessary for all the unconditionally valid syllogistic schemes in which the model achieves high downstream accuracy ($\geq$ 60%) (see Appendix A for the list of schemes). This result suggests that LMs learn reasoning mechanisms that are transferable across different schemes.
4. The intervention results on models with different architectures and sizes (i.e., GPT-2 (Radford et al., 2019), Pythia (Biderman et al., 2023), Llama (et al., 2024), Qwen (Yang et al., 2024)) show similar suppression mechanism patterns and information flow. However, we found evidence that the contribution of attention heads becomes more complex with model sizes, further supporting the hypothesis of increasing contamination from external world knowledge.
@@
Our study is conducted using TransformerLens (Nanda and Bloom, 2022) on an Nvidia A100 GPU with 80GB of memory. The dataset and code to reproduce our experiments are available online to encourage future work in the field [^1] https://github.com/neuro-symbolic-ai/Mechanistic-Interpretation-Syllogism.
@@ 2.2 Circuit Discovery
This technique involves modifying the activation of targeted components and observing the resulting changes. Our study primarily examines the activation of residual streams and multi-head self-attention at the sequence level to trace the term information flow. Activation patching includes three model runs (Clean, Corrupted and Intervened) alongside a quantification process to measure the effect of the interventions (see Figure 2(1)–(4)). Given a masked syllogistic input $x=[\mathcal{P}_{1};\mathcal{P}_{2};\mathcal{C}\setminus\{p\}]$ as $(s,m_{1},m_{2},p)$, which can be read as “All $s$ are $m_{1}$. All $m_{2}$ are $p$. Therefore, all $s$ are [mask]”, and its correct completion $y=p$, the activation patching technique consists of the following steps:
@@
For the target syllogistic input $(s,m_{1},m_{2},p)$, we record the baseline logit difference, $\delta_{+}(p,m)$ from the forward pass output of the model (Figure 2(1)).
@@
Figure 3: Visualization of the symbolic reasoning circuit identified in the model. (a) Attention output patching and (b) residual stream patching using middle-term interventions highlight the role of attention head $\mathbf{h}_{11.10}$ in aggregating information from the duplicated middle term to the predicate token position $[p]$. (c) Logit lens visualization of the OV (output vector) of $\mathbf{h}_{11.10}$ across 26 uppercase letters reveals its suppressive effect on the attended token at $[p]$. (d) Residual stream patching with all-term corruption indicates that key reasoning information is concentrated at term token positions. (e)–(h) Average attention weights across the batch: (e) attention from token position $[p]$, attending strongly to $[m_{2}]$; (f)–(h) attention from the last token position, identifying which heads (mover heads) propagate term information to the output. For clarity, a dash (–) on the axis indicates the averaged values for tokens appearing between terms.
2.3 Causal Interventions
@@ 3 Mechanistic Interpretation
We select the LM for circuit analysis based on a trade-off between model size and performance. To this end, we measure the accuracy of GPT-2 (Radford et al., 2019) at various sizes (117M, 345M, 762M, and 1.5B) on the syllogism completion task, aiming to identify potential phase transition points where performance changes significantly (Jason et al., 2022; Kaplan et al., 2020). We observe a marked transition from small to medium sizes, where average performance across three datasets increases by 439.06% (see Figure 6(a), Appendix E). The shift is even more pronounced under the logit difference metric (see Section 4.4 and Figure 6(b), Appendix E). Based on these findings, we select GPT-2 Medium—whose architecture is summarized in Appendix F—for circuit discovery, and subsequently evaluate generalization across model sizes and architectures (Sections 4.3 and 4.4).
@@ 3.1 Transitive Reasoning Mechanisms
We further investigate how information from $[m_{2}]$ is moved to $[p]$, positing that this relates to the model’s internal reasoning mechanism. We notice that at position $[p]$, the model can observe the complete AAA-1 syllogistic structure (Figure 2(a)) with middle-term duplication. We hypothesize that this structural information for reasoning is collected at one position, given that information refinement occurs at the specific position such as last token (Hanna et al., 2024; Stolfo et al., 2023). To verify this, we employ path patching (Wang et al., 2023), a more selective variant of activation patching, to trace the information flow of head $\mathbf{h}_{11.10}$. Our results show that $\mathbf{h}_{11.10}$ operates based on several induction heads (Elhage et al., 2021) ($\mathbf{h}_{5.8}$, $\mathbf{h}_{6.1}$, $\mathbf{h}_{6.15}$ and $\mathbf{h}_{7.2}$) formed at $[m_{2}]$. These heads attend to the $[[m_{1}]+1]$ token due to the $m_{1}\equiv m_{2}$ conditioned matching operation, likely containing $m_{1}$-related information from previous token heads. We conclude, therefore, that $m_{1}$ and $m_{2}$ information are aggregated at position $[p]$. Additional details of the path patching results are available in the Appendix G.
A suppressive mechanism is revealed through Logit Lens.
@@ 3.2 Term-Related Information Flow
Given the findings that the information for prediction resides in term-specific positions, we trace which attention heads transfer the information from each term position ($[p]$, $[m_{1}]$ and $[m_{2}]$) to the last token. We call these heads mover heads as the existence of “positive or negative mover heads” that carry or suppress information from the specific token position to the last token position (Wang et al., 2023; García-Carrasco et al., 2024). To identify the sources of information flow, we apply attention value-based patching and isolate nine notable mover heads (see Appendix H for localization details). Among them, $\mathbf{h}_{14.14}$, $\mathbf{h}_{15.14}$, and $\mathbf{h}_{18.12}$ act as positive copy heads, attending strongly to $[p]$, while $\mathbf{h}_{19.1}$ functions as a positive suppression head, exhibiting high attention to $[m_{2}]$ and $[s]$ (see Figure 3(f)). In contrast, $\mathbf{h}_{9.9}$, $\mathbf{h}_{11.1}$, $\mathbf{h}_{12.1}$, $\mathbf{h}_{17.2}$, and $\mathbf{h}_{23.10}$ serve as negative copy heads, focusing primarily on $[m_{1}]$ (Figure 3(g)) or showing more diffuse pattern (Figure 3(h)).
@@
3.3 Circuit Discovery Summary
Overall, we found that the mechanisms for syllogistic inference involve the following phases:
@@ 3.3 Circuit Discovery Summary
Token-specific information is propagated to the last position. The process ends in the prediction shift from ‘B’ to the correct token ‘C’.
These results suggest that the circuit is characterized by an internal error correction mechanism. Interestingly, this mechanism is different from the way human experts would reason on syllogistic arguments through the systematic application of abstract logical primitives and inference rules.
4 Circuit Evaluation
The circuit is sufficient and necessary for symbolic arguments.
@@ 4 Circuit Evaluation
Figure 4: Circuit ablation results for (a) correctness and (b) robustness on the symbolic dataset. Solid lines show numeric perturbation-based performance, dotted lines indicate quantifier perturbation-based performance in (b), and the dashed line shows the baseline logit difference without knockouts. These results validate the necessity, sufficiency, and robustness of the identified symbolic reasoning circuit.
Figure 5: Circuit ablation results for evaluating transferability on non-symbolic datasets: (a) the belief-consistent and (b) the belief-inconsistent. The dashed line represents the baseline logit difference without knockouts. In both cases, the symbolic circuit is necessary, as ablating it reduces performance. However, the circuit is only sufficient for belief-consistent data, as recovery fails in belief-inconsistent settings, suggesting interference from belief-bias.
The circuit is robust to superficial variations.
@@
4.1 Circuit Transferability to Concrete Arguments
The circuit is necessary for non-symbolic arguments, yet not sufficient for belief-inconsistent ones.
We present ablation results for the two generated non-symbolic datasets: belief-consistent (Figure 5(a)) and belief-inconsistent (Figure 5(b)). Notably, the symbolic circuit proves necessary for both types of non-symbolic inputs, suggesting that the logic derived from symbolic syllogisms remains essential even when natural words are substituted. Regarding sufficiency, while belief-consistent data show significant performance recovery, we observe an inability to restore performance on the belief-inconsistent set despite an increasing trend. These results indicate that belief biases encoded in different attention heads may play an important role.
@@ 4.2 Generalization to Syllogistic Schemes
Table 2: Generalizability across unconditionally valid syllogistic forms. Columns C1, C2, and C3 indicate whether conditions are met ($\checkmark$) or not (-): C1: Necessity, C2: Sufficiency, C3: Positive Logit Difference. ‘Accuracy (Acc)’ shows performance on the completion task for each syllogism. The results show that the model achieves high downstream accuracy particularly for affirmative syllogisms.
@@
We further aim to understand whether the symbolic circuit is specific to the AAA-1 syllogism (Barbara). To this end, we extend our experiment to encompass all 15 unconditionally valid syllogisms (see Appendix A and I for details about the syllogistic schemes and more experimental results). To evaluate the transferability to all 15 schemes, we verify three main conditions, reported in Table 2: (C1) Necessity, (C2) Sufficiency, and (C3) Positive Logit Difference. We also measure the accuracy of the completion task for each syllogism. From Table 2, we observe that the circuit is sufficient and necessary for all the syllogistic schemes in which the model achieves high downstream accuracy ($\geq$ 60%), supporting the conclusion that the circuit includes components that are crucial for the emergence of syllogistic reasoning in general. Notably, all three conditions are satisfied for the affirmative syllogisms (AII-3, IAI-3, IAI-4), for which the model achieves an accuracy above 0.6.
4.3 Generalization to Model Sizes